OpenStudy (anonymous):

determine the value of a so that the average rate of change of the function j(x)=3x²+ax-4 over the interval -1≤x≤2 is 8

OpenStudy (phi):

did you try to set up the equation ?

OpenStudy (anonymous):

i plugged the values of x into the equation and got two new equations if thats what you mean?

OpenStudy (anonymous):

i got f(2) = 10+ax f(-1)= 1+ax ... ?

OpenStudy (phi):

the equation is ( j(2) - j(-1))/(2 - -1) = 8

OpenStudy (tkhunny):

"average rate of change" Perhaps the "slope" between the endpoints?

OpenStudy (anonymous):

oh so i would just do ((10+ax)-(1+ax))/(2-(-1)) = 8?

OpenStudy (phi):

except j(2) should not have any x's in it (because x is replaced by 2) ditto for j(-1) so redo that part

OpenStudy (anonymous):

Oh wait ... so j(2) = 10+2a j(-1) = 1-1a ?

OpenStudy (phi):

yes

OpenStudy (anonymous):

and then what i said earlier right?

OpenStudy (phi):

though j(-1) is -1 - 1a , right ?

OpenStudy (anonymous):

\[\frac{ (10+2a)-(-1-1a) }{ 2-(-1) }\] right?

OpenStudy (anonymous):

wait... i think i did that wrong

OpenStudy (phi):

and j(2) with j(x)=3x²+ax-4 is 3*4 + 4a -4 = 12 +4a -4 = 4a+8

OpenStudy (anonymous):

\[\frac{ (10+2a)-(1-1a) }{ 2-(-1) }\]

OpenStudy (anonymous):

doesnt x=2 not 4?

OpenStudy (phi):

oops j(2) is and j(2) with j(x)=3x²+ax-4 is 3*2*2 + 2a -4 = 12 +2a -4 = 2a+8

OpenStudy (anonymous):

OH WOW!! IVE BEEN PUTTING 2 INSTEAD OF 4 for the end number

OpenStudy (phi):

and j(-1) is 3*-1*-1 + -1*a -4 = 3 -a -4 = -a - 1

OpenStudy (anonymous):

\[\frac{ 8+2a-(-1-a) }{ 2-(-1) }\]

OpenStudy (phi):

yes and that is equal to 8 (the average rate of change)

OpenStudy (anonymous):

soooooo \[\frac{ 9+3a }{ 3 }\]\[3+a\]\[a=8-3\]\[a=5\]

OpenStudy (anonymous):

yay ! thaankkkss @phi

OpenStudy (phi):

yw